or A yields products, time and concentration data were collected and plotted as shown here. [A](M): 0.800, 0.740, 0.680, 0.620 t(s): 0.0,30.0, 60.0, 90.0 Determine the reaction order, the rate constant, and the units of the rate constant. Order=? K= ? and units Comment
Answer #1

To determine the order, let's plot the data as if were the zero order reaction, first reaction order and then again for a second order reaction. The one with the straight slope (R2 closer to 1) will be the one with the correct order.

EDITED

For a zero order reaction, we plot A vs t, and the data obtained:
y = 0.8 - 0.002x
r2 = 1

This value of r2 = 1 means that the reaction has a very straight slope, so the reaction is order zero. However, let's see what happen in order 1 and 2.

For a first order reaction, we plot lnA vs t, and doing this, we obtain the following data:
y = -0.2196 - 0.00283x
r2 = 0.99856

This could be the order of reaction, but as we already see, zero order has a r = 1. but let's do it for the second order just to be sure. in this case, we plot 1/A vs t so:
y = 1.24 + 0.00403x
r2 = 0.9943

The closer value is the one with the zero order reaction, so, this is a zero order reaction. The slope will give us the rate constant k
k = 0.00283 s-1

Hope this helps